
KX Vel added the 30 s exposure requested in Note # 5.
TOI-2641 file 1129, tel moved ~ 5["] in dec (partially cloudy, low wind, HA ~ 2hr W.) Added an extra integration 360 [s] (file 1131) to 'compensate'.
DIMM down between 02:31 and 04:37 UT. Since 06:54 UT, down again (overcast)
File 1154, ThAr looks like a bias, renamed as 'junk'. Useful lamps, 1155 and 1156.
08:40 Too overcast, waiting.
Underline: next objects were NOT observed: 'T0290403522' (though tried), 'TIC273418421' (though tried), '139251123' and 'TIC303317324'. An awful night: in general, poor S/N ratio.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-06-16 | 9.0 | 0.0 | 0.0 | 3.0 | 0.0 | 12 |
Links
[1] mailto:thenry@astro.gsu.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39318/delete?destination=print/node/17850
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39318/edit?destination=print/node/17850
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/17850?destination=print/node/17850
None.